Problem solving .show your solution 1.An object is released from rest.How far does it fall during the 2nd second of its fall?


1
Expert's answer
2021-10-04T19:37:40-0400

Solving for d after 2 seconds

"d_2=1/2 gt^2=1/2\times 9.8\times 2^2=19.6m"

Solving for d after 1 second

"d_1=1/2gt^2=1/2\times 9.8\times 1^2=4.9m"

Solving for d during the 2nd second

d = d2 - d1

d = 19.6 m - 4.9 m

d = 14.7 m

The object has fallen 14.7 m during the second second of its fall.